# Wandering around: A bioinspired approach to visual attention through object motion sensitivity

### v1

@@ 3.2.1 Characterisation of the SNN Proto-Object model
RGB stimuli Event map OMS map Saliency map (from events) Saliency map (from OMS map)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: From left to right: RGB stimuli in the case of calibration circles, proto-object patterns, and non-proto-object patterns from the SalMapIROS dataset [iacono2019proto], Event map, OMS map, and Saliency map (the color legend follows the jet map, with blue indicating minimum values and red representing maximum values) with events and OMS map as input, respectively.
To validate the system’s functionality, we replicated the characterisation described in [iacono2019proto, ghosh2022event, Dangelo_etal22]. Events from the SalMapIROS dataset [iacono2019proto] were utilised to qualitatively evaluate the system’s performance across three primary scenarios: calibration circles (ideal stimuli), proto-object patterns, and non-proto-object patterns (see Figure 8). The proto-object pattern derives from the characterisation of the original RGB implementation [russell2014model] depicting corners and corners facing each other, which conceptually resembles a proto-object.
In [iacono2019proto] the events were collected using ATIS cameras mounted on the humanoid robot iCub [bartolozzi2011embedded] by generating small stereotypical circular eye movements, emulating fixational eye movements observed in many vertebrate species [martinez2008fixational]. During the experiments, the robot observed patterns placed on a desk (72 cm high), while the camera setup was positioned at a total height of approximately 124 cm (104 cm for the robot’s base height plus  20 cm for support). This range of sizes was selected based on the typical objects the robot can interact with, considering the constraints of its workspace and grasping capabilities [vezzani2017grasping, jamone2016benchmarking]. The characterisation analyses the responses feeding the SNN Proto-object model with raw events and the OMS map from the sOMS model.
In the calibration circles scenario, the system, with both inputs, detected almost four out of six circles, with radii ranging from 10 to 60 mm, corresponding to sizes between 25 and 50 pixels on the image plane. In the proto-object pattern, the model exhibited a strong response to corners facing each other representing proto-objects [russell2014model] while displaying a weaker response to corners in the non-proto-object pattern. Furthermore, the OMS map, due to Gaussian filtering, produces a thicker and clearer edge map, making it an ideal input candidate for the SNN Proto-object model. The parameters obtained from the characterisation are summarised in Table 1.
@@ 3.3 Spiking Attention Control
The Spiking Attention Control (sAC) obtains $P(x,y)$ from the SNN Proto-object model and generates the saccadic movements (pan-tilt) through a simple attention control model implemented using spiking neurons. The objective of the control system is to minimise the distance between the centroid of the most salient object and the center of the sensor’s field of view. This is achieved by moving the center of the DVS integrated into the Speck device mounted on the FLIR PTU to foveate towards the salient point (see Figure 1). To describe this feedback control system, we use the Neural Engineering Framework (NEF) [eliasmith2003neural]. Since the camera is mounted on a PTU, we can treat the axis-aligned errors and controls separately. A proportional controller is used, defined as:
@@
The neural controller was implemented using spiking LIF neurons within the Nengo software package [bekolay2014nengo]. The controller was instantiated by solving a least-squares minimization for the decoder weights, $\mathbf{w}$, of a neural network with a single hidden layer, assuming randomly generated and fixed input weights, $\mathbf{e}$. For the pan controller, the following equation is solved:
$\mathbf{w}_{\mathrm{pan}}=\underset{\mathbf{w}\in\mathbb{R}^{d}}{\arg\min}\|K_{\mathrm{pan}}\varepsilon_{x}-\mathbf{w}\cdot\left(h(t)\ast\mathbf{a}(\langle\mathbf{e}_{\mathrm{pan}},\varepsilon_{x}\rangle)\right)\|^{2}$ (8)
where $h(t)$ is a synapse model (a low-pass filter), $a(\cdot)$ is the activity for the $d$ leaky integrate and fire neurons (bias term omitted for compactness) in the population representing the pan error, $\mathbf{e}_{\mathrm{pan}}$ are the randomly generated encoding weights for the pan controller, and $\varepsilon_{x}=x_{\mathrm{obj}}-x_{\mathrm{center}}$, is the control error given the current saliency input. The same technique is used to solve for the tilt control weights, $\mathbf{w}_{\mathrm{tilt}}$. We exploit Kirchoff’s current laws to compute the difference between the actual and desired location of the saliency maximum directly in the current injected into the neurons. For a fixed target location, this difference could also be encoded in the biases of the neurons that make up the controller.
To avoid interference between the two control dimensions, a separate neural population is maintained for both the pan and tilt commands. Proportional control gains were hand-tuned to produce the desired performance. The final values of $u_{\text{pan}}$ and $u_{\text{tilt}}$ are obtained from the $cmd_{pan}$ and $cmd_{tilt}$ output commands of the sAC control as follows:
$u_{\text{pan}}=\left\lfloor\frac{\text{cmd}_{\text{pan}}}{2\cdot\alpha\cdot\text{degrees\_per\_pos}}\right\rfloor$ (9)
[…]
@@ 3.3.1 Characterisation of the Spiking Attention Control
To evaluate the accuracy of the sAC model in foveating towards a desired point, 10 trials were conducted, and the displacement in $x$ and $y$ pixel coordinates from the sAC pixels commands, $cmd_{pan}$ and $cmd_{tilt}$, to the salient point $P(x,y)$, were analysed. The results showed a mean displacement of $11.76$ pixels (standard deviation: $2.58$) in the vertical ($y$) direction and $4.64$ pixels (standard deviation: $1.49$) in the horizontal ($x$) direction, indicating greater accuracy in the horizontal axis.
@@ 3.4 Real-world demonstrator: Bioinspired attention system
A reference processing architecture (Figure 1) has been implemented with the aim to streamline the data flow and maximise computational efficiency. Some optimisations are crucial for handling high-frequency event data, which is often encountered in real-time applications such as robotic vision systems. Parallel processing for pan-tilt movement management, noise reduction and minimising redundant computation enable real-time operation with high responsiveness and accuracy, even for large volumes of event data. Below, we outline the key aspects of our reference implementation demonstrating the operation of the proposed event-based visual attention pipeline [^5] The code can be found at https://github.com/GiuliaDAngelo/Speckegomotion/tree/demo.
The pipeline consists of several initialisation steps followed by a loop that processes events arriving from the event source, which is the DVS camera integrated into the Speck device. The object motion sensitivity module, the proto-object detection module, the visual attention module with the Nengo controller and the PTU are initialised with the parameters outlined in Table 1. Additionally, the number of neurons ($n$) in the system’s convolutional layers is directly related to the number of units in the output feature map, which is determined by the input size, kernel size, and stride.
@@
$\alpha$ = 0.80 $\omega$ = 3 $k_{pan}$=$k_{tilt}$=1
@@
Orientations: [$0\degree$, $45\degree$, $90\degree$, $135\degree$]
@@
The reference implementation focuses on demonstrating an end-to-end bioinspired attention system that uses event-based data and pan-tilt movements to foveate the most salient point in the visual field, performing fixational eye movements to detect the next point, as shown in Figure 1. To achieve this, the pipeline implements a loop that accumulates events over a short period of time (2 ms) and feeds them into the object motion sensitivity module, which outputs and highlights regions of non-coherent motion.
@@
[…]
@@
The same applies for the $\text{tilt}_{range}$ since the resolution of the DVS is 128$\times$128. Finally, the pan and tilt ranges are clipped to ensure they remain within the physical limits of the PTU.
Once the PTU has directed the fovea of the system towards the salient point, the PTU performs the fixational eye movements [ko2010microsaccades]. The fixational eye movements are simulated using a random walk mechanism. For each eye movement $i$ in a sequence of $N$ fixational eye movements, the pan and tilt angles are updated by adding a random perturbation $r_{i}$:
for $\text{for }i\in\{1,2,\dots,N\}:$
[…]
These updated angles are sent as commands to the PTU, simulating the small, involuntary movements characteristic of fixational eye movements and necessary to process the new information of the visual field and detect the next salient point.
@@
Initialise $OMS\_module$, $attn\_module$ and $Nengo\_controller$
@@
       issue blocking command to PTU unit to move by $u\_pan$, $u\_tilt$
@@ 4 Experiments & Results
The experiments, both qualitative and quantitative, evaluate the system as a whole, quantifying the advantages of utilising the sOMS model for object motion segmentation and enhancing multi-object detection through the SNN Proto-object model. The sOMS was initially evaluated to quantify the proportion of events suppressed that are unlikely to correspond to moving objects. To assess the model’s accuracy in effectively segmenting multiple objects, it was benchmarked on the EVIMO [mitrokhin2019ev] dataset, measuring mean IoU and mean SSIM percentages. To further evaluate the contribution of the SNN Proto-object model, we qualitatively compared results with and without its inclusion and quantitatively assessed its accuracy in multi-object detection. To further leverage the inherent advantages of event-driven cameras, the sOMS model and the SNN Proto-object model have been qualitatively and quantitatively evaluated on the LLE-VOS [li2024event] dataset under challenging indoor and outdoor lighting conditions to assess the accuracy of the SNN Proto-object model in object detection. Finally, the entire closed-loop bioinspired attention system has been demonstrated, with the Speck device mounted on the PTU successfully performing saccades toward the salient point of interest and generating fixational eye movements to detect the next salient point (see the accompanying video [^6] https://youtu.be/dcAJlDgVR0o).
Figure 9: An example from one of the 10 trials illustrating the event counts before and after applying the sOMS model to events received from the DVS sensor (128$\times$128 pixels). The events count from the DVS is represented in red and the outcoming events from the sOMS model in blue. The y-axis shows the event count, while the x-axis represents time in seconds.
@@
As described in Section 3.1.1, the sOMS model enhances contrast in motion, enabling objects with relative motion with respect to the background to emerge. Conversely, this leads to the exclusion of a substantial number of events unrelated to object motion, such as those originating from the static background, thereby reducing the processing load. This demonstrates an additional advantage of the proposed model. To quantitatively validate this observation, the number of events was recorded every 2 ms before and after the sOMS model (see Table 2).
@@
To validate the ability of the sOMS model (as described in Section 3.1) to enhance and segment object motion, the model’s response was benchmarked using the EVIMO dataset [mitrokhin2019ev]. The EVIMO dataset [mitrokhin2019ev] is a motion segmentation dataset designed for event cameras. It provides high-resolution temporal data for egomotion estimation and motion segmentation, featuring pixel-wise object masks and 6-DoF poses for both the camera and the moving objects.
Sub-dataset Event map [mitrokhin2019ev] Ground Truth [mitrokhin2019ev] OMS map mean IoU % [clerico2024retina] mean IoU % mean SSIM %
Box Refer to caption Refer to caption Refer to caption 72 $\pm$ 16 64.79 $\pm$ 0.02 89 $\pm$ 0.08
Fast Refer to caption Refer to caption Refer to caption 69 $\pm$ 3 69.85 $\pm$ 0.15 90 $\pm$ 0.06
Floor Refer to caption Refer to caption Refer to caption 94 63.21 $\pm$ 0.22 94$\pm$ 0.22
Table Refer to caption Refer to caption Refer to caption 88 $\pm$ 10 73.59 $\pm$ 0.22 89 $\pm$ 0.11
Tabletop Refer to caption Refer to caption Refer to caption 72 $\pm$ 14 82.24 $\pm$ 0.18 96 $\pm$ 0.06
Wall Refer to caption Refer to caption Refer to caption 82±6 64.49 $\pm$ 0.07 84 $\pm$ 0.04
Figure 10: From left to right: sub-dataset name, the Event map, ground truth binary mask, OMS map and mean IoU percentage for the six EVIMO [mitrokhin2019ev] sub-datasets.
The sOMS model for the benchmark employs $\alpha=0.80$ and $\tau=0.02$ to align with the EVIMO dataset’s Event map and mask collection. In Figure 10, the model is qualitatively and quantitatively assessed by comparing the pixel-wise object masks with the OMS maps across six different sub-datasets: Box, Fast, Floor, Table, Tabletop, and Wall. Each sub-dataset contains different sequences of motion: Box (6), Fast (3), Floor (2), Table (4), Tabletop (4), and Wall (2). To quantitatively assess the sOMS model, the mean IoU percentage (as described in [clerico2024retina]) and the SSIM were computed between the ground truth mask and the OMS map (Equation 10).
@@
Figure 10 presents the benchmark results against [clerico2024retina], demonstrating comparable performance in terms of IoU percentage for the Fast dataset, lower performance for the Box, Table, and Wall datasets, and improved performance for the Tabletop dataset. The SSIM results indicate a superior and more consistent performance of the model compared to the insights provided by the IoU values. The observed difference in performance between SSIM and IoU can be attributed to the fact that SSIM focuses on capturing structural similarities and subtle texture variations, making it more forgiving of small discrepancies. In contrast, IoU is more rigid, disproportionately penalizing even minor misalignments or boundary errors in object overlap. Furthermore, the results from the sOMS model are slightly lower than those of the OMS model by Clerico et al. [clerico2024retina], which can be attributed to the sparse response characteristics of the spiking implementation.

### v2

@@ 3.2.1 Characterisation of the SNN Proto-Object model
Figure 8: From left to right: RGB stimuli in the case of calibration circles, proto-object patterns, and non-proto-object patterns from the SalMapIROS dataset [30], Event map, OMS map, and Saliency map (the color legend follows the jet map, with blue indicating minimum values and red representing maximum values) with events and OMS map as input, respectively.
To validate the system’s functionality, we replicated the characterisation described in [30, 31, 39]. Events from the SalMapIROS dataset [30] were utilised to qualitatively evaluate the system’s performance across three primary scenarios: calibration circles (ideal stimuli), proto-object patterns, and non-proto-object patterns (see Figure 8). The proto-object pattern derives from the characterisation of the original RGB implementation [15] depicting corners and corners facing each other, which conceptually resembles a proto-object.
In [30] the events were collected using ATIS cameras mounted on the humanoid robot iCub [29] by generating small stereotypical circular eye movements, emulating fixational eye movements observed in many vertebrate species [84]. During the experiments, the robot observed patterns placed on a desk (72 cm high), while the camera setup was positioned at a total height of approximately 124 cm (104 cm for the robot’s base height plus  20 cm for support). This range of sizes was selected based on the typical objects the robot can interact with, considering the constraints of its workspace and grasping capabilities [85, 86]. The characterisation analyses the responses feeding the SNN Proto-object model with raw events and the OMS map from the sOMS model.
In the calibration circles scenario, the system, with both inputs, detected almost four out of six circles, with radii ranging from 10 to 60 mm, corresponding to sizes between 25 and 50 pixels on the image plane. In the proto-object pattern, the model exhibited a strong response to corners facing each other representing proto-objects [15] while displaying a weaker response to corners in the non-proto-object pattern. Furthermore, the OMS map, due to Gaussian filtering, produces a thicker and clearer edge map, making it an ideal input candidate for the SNN Proto-object model. The parameters obtained from the characterisation are summarised in Table 1.
@@ 3.3 Spiking Attention Control
The Spiking Attention Control (sAC) obtains $P(x,y)$ from the SNN Proto-object model and generates the saccadic movements (pan-tilt) through a simple attention control model implemented using spiking neurons. The objective of the control system is to minimise the distance between the centroid of the most salient object and the center of the sensor’s field of view. This is achieved by moving the center of the DVS integrated into the Speck device mounted on the FLIR PTU to foveate towards the salient point (see Figure 1). To describe this feedback control system, we use the Neural Engineering Framework (NEF) [87]. Since the camera is mounted on a PTU, we can treat the axis-aligned errors and controls separately. A proportional controller is used, defined as:
@@
The neural controller was implemented using spiking LIF neurons within the Nengo software package [81] and was instantiated by solving a least-squares optimization problem to determine the decoder weights, $\mathbf{w}$, of a single-hidden-layer neural network, assuming fixed, randomly generated input weights, $\mathbf{e}$. For the pan controller, the following equation is solved:
$\mathbf{w}_{\mathrm{pan}}=\underset{\mathbf{w}\in\mathbb{R}^{d}}{\arg\min}\left\|\frac{K_{\mathrm{pan}}}{\text{degrees\_per\_pos}}\varepsilon_{x}-\mathbf{w}\cdot\left(h(t)\ast\mathbf{a}(\langle\mathbf{e}_{\mathrm{pan}},\varepsilon_{x}\rangle)\right)\right\|^{2}$ (7)
where $h(t)$ is a synapse model (a low-pass filter), $a(\cdot)$ is the activity for the $d$ leaky integrate and fire neurons (bias term omitted for compactness) in the population representing the pan error, $\mathbf{e}_{\mathrm{pan}}$ are the randomly generated encoding weights for the pan controller, and $\varepsilon_{x}=x_{\mathrm{obj}}-x_{\mathrm{center}}$, is the control error given the current saliency input. Here we make explicit the term degrees_per_pos, which is the angular resolution. The same technique is used to solve for the tilt control weights, $\mathbf{w}_{\mathrm{tilt}}$. We exploit Kirchoff’s current laws to compute the difference between the actual and desired location of the saliency maximum directly in the current injected into the neurons. For a fixed target location, this difference could also be encoded in the biases of the neurons that make up the controller.
To avoid interference between the two control dimensions, a separate neural population is maintained for both the pan and tilt commands. Proportional control gains were hand-tuned to produce the desired performance. The final values of $cmd_{\text{pan}}$ and $cmd_{\text{tilt}}$ are obtained from the $u_{\text{pan}}$ and $u_{\text{tilt}}$ output commands of the sAC control as follows:
$\text{cmd}_{\text{pan}}=\left\lfloor u_{\text{pan}}\right\rfloor,$ (8)
where $\lfloor\cdot\rfloor$ represents the floor operator. The same relation applies for $u_{tilt}$. While this work employs a simple proportional controller to direct eye movements, these neural control techniques can be extended to more complex feedback controllers, including adaptive components (e.g.[88, 89]), although this is beyond the scope of the current work.
@@ 3.3.1 Characterisation of the Spiking Attention Control
To evaluate the accuracy of the sAC model in foveating towards a desired point, 10 trials were conducted, and the displacement in $x$ and $y$ pixel coordinates from the sAC pixels commands, $cmd_{\text{pan}}\cdot\text{degrees\_per\_pos}$ and $cmd_{\text{tilt}}\cdot\text{degrees\_per\_pos}$, to the salient point $P(x,y)$, were analysed. The results showed a mean displacement of $11.76$ pixels (standard deviation: $2.58$) in the vertical ($y$) direction and $4.64$ pixels (standard deviation: $1.49$) in the horizontal ($x$) direction, indicating greater accuracy in the horizontal axis.
@@ 3.4 Real-world demonstrator: Bioinspired attention system
The real-world implementation has been developed (Figure 1) to streamline the data flow and maximise computational efficiency. Some optimisations are crucial for handling high-frequency event data, commonly encountered in real-time applications such as robotic vision. Parallel processing for pan-tilt movement management, noise reduction, and the minimisation of redundant computations enables responsive and accurate real-time performance, even with large volumes of event data. Below, we outline the key aspects of our implementation, demonstrating the operation of the proposed event-based visual attention pipeline [^5] The code can be found at https://github.com/GiuliaDAngelo/Speckegomotion/tree/demo.
The pipeline consists of several initialisation steps followed by a loop that processes events arriving from the event source. The object motion sensitivity module, the proto-object detection module, the visual attention module with the Nengo controller and the PTU are initialised with the parameters outlined in Table 1. Additionally, the number of neurons ($n$) in the system’s convolutional layers is directly related to the number of units in the output feature map, which is determined by the input size, kernel size, and stride.
@@
$\alpha$ = 0.80 $\omega$ = 3 $k_{pan}$$k_{tilt}$=1
@@
Orientations: $0\degree$, $45\degree$, $90\degree$, $135\degree$]
@@
The implementation focuses on demonstrating an end-to-end bioinspired attention system that uses event-based data and pan-tilt movements to foveate the most salient point in the visual field, performing fixational eye movements to detect the next point, as shown in Figure 1. To achieve this, the pipeline implements a loop that accumulates events over a short period of time (2 ms) and feeds them into the object motion sensitivity module, which outputs and highlights regions of non-coherent motion.
@@
The pan and tilt ranges are defined based on the DVS angular resolution, field of view, and physical constraints of the PTU. The angular resolution is determined by the conversion factor $\text{degrees\_per\_pos}=0.02572^{\circ}$, which corresponds to one PTU position of $92.5714$ arcseconds ($92.5714\,\text{arcsec}\times 0.0002778^{\circ}/\text{arcsec}=0.02572^{\circ}$). The field of view is calculated using the focal length of the DVS sensor $f=1.7\,\text{mm}$ and the horizontal sensor size $s_{\text{hor}}=5.12\,\text{mm}$. The horizontal field of view $\text{FOV}_{\text{hor}}$ is derived as:
@@
The same applies for the $\text{tilt}_{range}$ since the resolution of the DVS is $\times128$. Finally, the pan and tilt ranges are clipped to ensure they remain within the physical limits of the PTU.
Once the PTU has oriented the system’s fovea toward the salient point, it initiates a sequence of fixational eye movements, mirroring those observed in biological vision systems [58]. These fixational eye movements are modelled using a random walk mechanism. Specifically, for each movement $i$ in a sequence of $N$ movements, both the pan and tilt angles are updated by adding a small random perturbation $r_{i}$, drawn from a uniform distribution:
$\text{for }i\in\{1,2,\dots,N\}:$
$\text{pan\_angle}\leftarrow\text{pan\_angle}+r_{i},\quad\text{tilt\_angle}\leftarrow\text{tilt\_angle}+r_{i}.$
Here, $N=8$ and the perturbation range is [60,60][-60,60] degrees. This choice of parameters was empirically determined to effectively generate events from the static scene while providing sufficient time for spatiotemporal integration of the information. These updated angles are sent as commands to the PTU, simulating the small, involuntary movements characteristic of fixational eye movements and necessary to process the new information of the visual field and detect the next salient point.
@@
Initialise OMS_moduleOMS\_module, attn_moduleattn\_module and Nengo_controllerNengo\_controller
@@
   issue blocking command to PTU unit to move by u_panu\_pan, u_tiltu\_tilt
@@ 4 Experiments & Results
The experiments, both qualitative and quantitative, evaluate the system as a whole, quantifying the advantages of utilising the sOMS model for object motion segmentation and enhancing multi-object detection through the SNN Proto-object model. The sOMS was initially evaluated to quantify the proportion of events suppressed that are unlikely to correspond to moving objects. To assess the model’s accuracy in effectively segmenting multiple objects, it was benchmarked on the EVIMO [59] dataset, measuring mean IoU and mean SSIM percentages. To further evaluate the contribution of the SNN Proto-object model, we qualitatively compared results with and without its inclusion and quantitatively assessed its accuracy in multi-object detection. To further leverage the inherent advantages of event-driven cameras, the sOMS model and the SNN Proto-object model have been qualitatively and quantitatively evaluated on the LLE-VOS [60] dataset under challenging indoor and outdoor lighting conditions, where we calculated the accuracy of the SNN Proto-object model in object detection. Finally, the entire closed-loop bioinspired attention system has been demonstrated, with the Speck device mounted on the PTU successfully performing saccades toward the salient point of interest and generating fixational eye movements to detect the next salient point (see the accompanying video [^6] https://youtu.be/dcAJlDgVR0o).
Figure 9: An example from one of the 10 trials illustrating the event counts before and after applying the sOMS model to events received from the DVS sensor ($\times128$ pixels). The events count from the DVS is represented in red and the outcoming events from the sOMS model in blue. The y-axis shows the event count, while the x-axis represents time in seconds.
@@
As described in Section 3.1.1, the sOMS model enhances motion contrast, allowing objects with relative motion to the background to emerge more distinctly. Simultaneously, it suppresses a substantial number of events unrelated to object motion, such as those caused by static background elements or sensor noise. Unlike generic event-reduction techniques that indiscriminately discard data, the sOMS model reduces events not by simply lowering the count (event-based cameras can potentially generate tens of millions of events per second) but by achieving targeted suppression based on motion relevance. This preserves salient object features while improving the efficiency of downstream processing. As a result, the computational load is significantly reduced without compromising task-relevant information, highlighting a key advantage of the proposed model. To quantitatively validate this observation, the number of events was recorded every 2 ms before and after the application of the sOMS model (see Table 2).
@@
To validate the ability of the sOMS model (as described in Section 3.1) to enhance and segment object motion, the model’s response was benchmarked using the EVIMO dataset [59]. The EVIMO dataset [59] is a motion segmentation dataset designed for event cameras. It provides high-resolution temporal data for egomotion estimation and motion segmentation, featuring pixel-wise object masks and 6-DoF poses for both the camera and the moving objects.
Figure 10: From left to right: Sub-dataset name, Event map, Ground truth binary mask, OMS map, mean IoU percentage and mean SSIM across the six EVIMO [59] sub-datasets. The figure presents results from [52] alongside our own for comparison.
The sOMS model for the benchmark employs $\alpha=0.80$ and $\tau=0.02$ to align with the EVIMO dataset’s Event map and mask collection. In Figure 10, the model is qualitatively and quantitatively assessed by comparing the pixel-wise object masks with the OMS maps across six different sub-datasets: Box, Fast, Floor, Table, Tabletop, and Wall. Each sub-dataset contains different sequences of motion: Box (6), Fast (3), Floor (2), Table (4), Tabletop (4), and Wall (2). To quantitatively assess the sOMS model, the mean IoU percentage (as described in [52]) and the SSIM were computed between the ground truth mask and the OMS map (Equation 9).
@@
Figure 10 presents the benchmark results against [52], demonstrating comparable performance in terms of IoU percentage for the Fast dataset, lower performance for the Box, Table, and Wall datasets, and improved performance for the Tabletop dataset. The SSIM results indicate a superior and more consistent performance of the model compared to the insights provided by the IoU values. The observed difference in performance between SSIM and IoU can be attributed to the fact that SSIM focuses on capturing structural similarities and subtle texture variations, making it more forgiving of small discrepancies. In contrast, IoU is more rigid, disproportionately penalizing even minor misalignments or boundary errors in object overlap. Furthermore, the results from the sOMS model are slightly lower than those of the OMS model by Clerico et al. [52], which can be attributed to the sparse response characteristics of the spiking implementation.